### v2

# TableCenterNet: A one-stage network for table structure recognition

Abstract
Table structure recognition aims to parse tables in unstructured data into machine-understandable formats. Recent methods address this problem through a two-stage process or optimized one-stage approaches. However, these methods either require multiple networks to be serially trained and perform more time-consuming sequential decoding, or rely on complex post-processing algorithms to parse the logical structure of tables. They struggle to balance cross-scenario adaptability, robustness, and computational efficiency. In this paper, we propose a one-stage end-to-end table structure parsing network called TableCenterNet. This network unifies the prediction of table spatial and logical structure into a parallel regression task for the first time, and implicitly learns the spatial-logical location mapping laws of cells through a synergistic architecture of shared feature extraction layers and task-specific decoding. Compared with two-stage methods, our method is easier to train and faster to infer. Experiments on benchmark datasets show that TableCenterNet can effectively parse table structures in diverse scenarios and achieve state-of-the-art performance on the TableGraph-24k dataset. Code is available at https://github.com/dreamy-xay/TableCenterNet.
Keywords Document Understanding Table Structure Recognition One-stage End-to-End
1 Introduction
As the main carrier of structured data, tables are prevalent in diverse media such as financial statements, academic papers, and natural-scene images, carrying the organization and delivery of key information. With the deepening digital transformation, the demand for accurately parsing the machine-understandable formats (including the spatial and logical locations of cells) of tables from unstructured data is constantly growing. Although humans can easily understand tables of various styles and layouts, it remains challenging for automated systems.
Previously, most table structure recognition (TSR) methods based on deep learning adopted a two-stage processing flow. For example, studies such as [1, 2, 3] first located the table region through object detection and then performed structural parsing on the cropped region, as illustrated in 1(a). Although these methods can handle some standard document tables, they are difficult to cope with the problem of geometric deformation in natural scenes due to shooting angle, occlusion or bending. In recent years, some works have attempted to optimize the two-stage paradigm. For instance, TGRNet [4] and LORE [5] sequentially predicted the spatial location and logical indices of cells through two networks, as shown in 1(b). While these methods perform well in different scenarios, their phased design requires independent training of multiple sub-modules to optimize performance, which significantly increases the training complexity. Furthermore, the serial inference process introduces additional computational overhead, making it difficult to meet the real-time requirements. Although the existing one-stage methods simplify the process, they face bottlenecks in scenario generalization. For example, TableNet [6] rely on the assumption of a clean and structured background and cannot handle blurring, uneven illumination, or geometric deformation in natural scenes. Although Cycle-CenterNet [7] and SCAN [8] support table parsing in natural scenes, it does not consider borderless tables and has insufficient accuracy in complex scenarios such as cross-row and cross-column merging. TRACE [9] is sensitive to geometric distortions such as bending and relies on complex post-processing algorithms. Overall, existing methods struggle to achieve an effective balance between cross-scenario adaptability, structural recognition robustness, and computational efficiency.
To address the above problems, we propose a one-stage end-to-end table structure parsing network, TableCenterNet, based on the CenterNet [10] framework. As illustrated in 1(c), this method unifies the spatial location detection and logical location prediction of cells into a parallel regression task for the first time, and realizes efficient joint learning through the synergistic architecture of a shared feature extraction layer and task-specific decoding. This design not only avoids the error accumulation problem of the two-stage approach, but also significantly improves the parsing accuracy of complex tables (e.g., merging cells across rows and columns) by implicitly learning the spatial-logical location mapping law. Experiments show that TableCenterNet is highly competitive and outperforms existing state-of-the-art methods for logical location prediction on the TableGraph-24k [4] dataset.
The main contributions of this paper are as follows:
We propose an end-to-end table structure parsing network that simultaneously predicts the spatial and logical locations of cells through parallel regression.
Experiments on three public benchmark datasets containing tabular images of diverse scenarios demonstrate the effectiveness of our proposed method.
We provide a simple and effective one-stage TSR method that simplifies the training process and accelerates inference, and the code is available to support further research on TSR.
2 Related Work
Early TSR methods [11, 12, 13, 14, 15, 16] primarily rely on manually designed features and heuristic rules. For example, they parse tables using visual cues such as text arrangement, lines, and templates. These methods are effective only for tables in specific formats (e.g., PDF) and struggle to handle complex tables. With the rise of statistical machine learning, approaches such as [17, 18, 19, 20] attempt to reduce the reliance on rules, but still require manual feature design, the strong assumptions regarding table layouts restrict their practical applications. In recent years, deep learning-based methods have shown better performance than traditional methods [21]. Based on the descriptions of these methods, we can broadly categorize them into three groups: row and column division-based methods, markup sequence generation-based methods, and cell reconstruction-based methods.
2.1 Row and column division-based methods
These methods mainly reconstruct the table structure by detecting or segmenting row and column regions [1, 6, 22, 23, 24, 25, 26, 3], or by directly predicting the separation lines for rows and columns [27, 28, 29, 30, 31, 32, 9]. For example, DeepTabStR [22] improves the object detection model using deformable convolutions to directly locate rows and columns; DeepdeSRT [1] and TableNet [6] utilize semantic segmentation models to predict row and column regions and generate cells through intersection operations. However, none of the above methods can handle the cases of row-spanning and column-spanning. Therefore, SPLERGE [27] further proposes a two-stage strategy of segmentation and merging, using a segmentation model to generate an initial cell grid and then processing cross-cells through a merging module. But these methods rely on the strict alignment assumption of rows and columns and have poor adaptability to curved tables or non-axis-aligned layouts. Although RobusTabNet [32] enhances robustness to distorted tables by introducing a spatial CNN module [33], the heuristic rules in its post-processing stage may still affect the overall performance due to low-quality separator masks.
2.2 Markup sequence generation-based methods
Such methods [34, 35, 36, 37, 38] treat TSR as a generation task from images to tag sequences and directly output structured tags such as HTML or LaTeX. TableMaster [37] is built on the Transformer architecture, which extracts visual features through the encoder and generates HTML sequences via the decoder. TableFormer [38] not only decodes structural tags but also decodes cell coordinates. However, these methods usually rely on large-scale training data, and the length of the generated markup sequence increases with greater table complexity, resulting in performance degradation. In addition, since the model uses a sequential decoding process, its inference is also more time-consuming.
2.3 Cell reconstruction-based methods
Cell reconstruction-based methods [39, 40, 41, 4, 42, 7, 43, 44, 45, 8, 5] usually predicts table cells first and then utilize these cells as the basic units to construct the global table structure by analyzing the associative information between them. For example, GraphTSR [39] models the predicted cells as graph nodes and employs a graph attention network to predict horizontal, vertical, or irrelevant relationships. FLAG-Net [41] enhances the reasoning ability for cell associations through both dense and sparse context modeling. TGRNet [4] and LORE [5] perform logical location prediction on the detected cells using Graph Convolutional Networks and Cascade Transformers, respectively, and directly generate logical indices. Existing methods generally adopt a two-stage process of "detection-association". Our method, however, makes full use of the cell location distribution information in the feature map during the detection stage, performs parallel regression of spatial and logical location, and directly matches the physical coordinates with the logical locations through interpolation maps to obtain the logical indices, thus simplifying the two-stage process into a one-stage end-to-end parsing.
3 Methodology
3.1 Overall Architecture
The overview of the proposed one-stage TableCenterNet is shown in Fig. 2. It employs a CNN backbone to extract visual features of cells from the input image, and then jointly predicts the spatial locations and logical locations of cells in a multi-task manner through six parallel regression heads. is the number of cells, and the definitions of and are as follows:
Where denotes the coordinates of the -th corner point of the -th cell. The four corner points are arranged clockwise starting from the upper-left corner. , , , and correspond to the four logical indices of the -th cell: starting-row, ending-row, starting-column, and ending-column, respectively.
Specifically, TableCenterNet comprises three main components: key point detection, spatial location regression, and logical location regression. The key point detection aims to streamline the joint prediction of spatial and logical location. Similar to CenterNet [10], we use two independent heads to regress the offsets and heatmaps of key points. The two channels of the key point heatmap correspond to the predictions of the cell center point and the corner coordinates, respectively. For the heatmap of each channel, when , it indicates the position of a key point, while signifies the background area, where and denote the height and width of the input image.
To regress the spatial locations of cells, we again incorporate two regression heads: one to predict the vectors from the cell centers to the corner points, and the other to predict the vectors from the corner points back to the centers. The last two regression heads are used to predict the row and column span information of the cells and to generate the logical location interpolation map. All regression results will be parsed and processed sequentially, and finally the table structure information will be output.
3.2 Spatial Location Regression
To enhance the accuracy of spatial location prediction, we adopted the Cycle-Pairing Module [7]. By regressing the vectors from the center of each cell to its four corners (center-to-corners) and the vectors from the four corners to the center (corners-to-center), and based on the center coordinates of each cell, we can calculate the approximate spatial locations of the cells. The specific calculation formula is as follows:
In addition, we have optimized PairLoss [7] to work more effectively in conjunction with logical position regression, while reducing the interference caused by invalid corners-to-center vectors. The formula for the improved spatial location regression loss is as follows:
Among them, and represent the ground truth (GT) corresponding to the predicted vectors and , respectively. denotes the mean absolute error [46] loss function, while represents the loss of invalid corners-to-center vectors. The hyperparameters , , and are used to adjust the relative importance of these loss items. Additionally, dynamically weights the loss based on the quality of the regression, it is defined as follows:
Using the corners as reference points, each corner can point to the centers of up to four different cells, resulting in corners-to-center vectors obtained through regression, where denotes the total number of cell corners. Since , in addition to the vectors matched by cells, there are also many invalid corners-to-center vectors, which are defined as . From this, the loss is formulated as follows:
3.3 Logical Location Regression
Considering that the extracted feature maps already contain the location information of cells, we capture the row-column demarcation features of the table by regressing the corresponding logical location interpolation map, obtaining the row interpolation map and the column interpolation map . Meanwhile, based on the cell center, we regress the row-column span of each cell, with the purpose of synergistically supervise the regression of the interpolation map and simplify the parsing process. Among them, and denote the row and column span of the -th cell respectively.
Generation of interpolation maps. Before training, we generate corresponding interpolation maps for each table image. For the GT of each image, we pair the logical indices of each cell with the physical coordinates to obtain two sets of polygons and for row and column interpolation. Where and can be defined as follows:
Then, we input and into Alg. 1 to calculate the corresponding row interpolation map and column interpolation map . Since the physical coordinates of the input cells are the same in the two calculations, the obtained mask image is also the same. The visualization results are shown in Fig. 3. During training, we only need to reduce , , and by a factor of 4 to use them as the GT for interpolation map regression.
Supervision of cell boundaries and spans. When performing logical location regression, two supervision methods are proposed to better understand the correspondence between logical and spatial locations and the constraints between logical locations. Supervision of cell boundary requires that all logical boundaries of cells can effectively anchor the row-column demarcation of the table and ensure the mutual exclusivity of the logical locations of different cells. Specifically, the loss scheme for cell boundaries can be expressed as follows:
Where and are weight functions used to enhance the supervision of the row-column demarcation of the table. In the interpolation map, the closer to the row-column demarcation, the higher the confidence of logical indices, and thus the weight should increase accordingly. and can be defined as follows:
Supervision of cell spans stipulates that the logical indices of each cell in the interpolation map should be consistent with its spans. Formally, the loss for constraining cell spans can be defined as follows:
Among them, the loss of spanning regression and the loss of spanning supervision for logical location of the i-th cell can be described by the following formulas:
In formula (12), denotes the GT corresponding to , and are weight functions used to calculate the average distance between and , which helps to coordinate the regression of the interpolation map with the cell span and accelerates the convergence of the model. and can be formulated as follows:
Where the corresponding row span and column span of the i-th cell in the interpolation maps and can be calculated by the following formulas:
Then the regression loss of the logical location are as:
3.4 Parsing Processing
In 4(a), obtained through regression are only approximations. The detected cell corner points (see 4(b)) need to be used for alignment processing [7] to obtain more accurate spatial locations , as shown in 4(c). Then, calculate the physical coordinates of all cells through 4x magnification.
Subsequently, by directly matching the physical coordinates with the interpolation maps, the logical indices of all cells can be obtained. The matching process for the logical indices of each cell can be formulated as follows:
Where denotes the coordinates of the upper-left corner of the physical coordinates of the i-th cell.
3.5 Training Strategy
Design of overall loss function. The proposed TableCenterNet is trained in an end-to-end manner for three optimization tasks: keypoint detection, spatial location regression and logical location regression. The global optimization can be defined as:
Where is the keypoint detection loss, which mainly consists of the heat map loss and offset loss of keypoints, consistent with CenterNet [10].
Selective gridding interpolation. For datasets with unbiased prior knowledge of physical and logical structures and no significant deformation, there is no need to generate interpolation maps based on cells before training. As shown in Fig. 5, the cells can be converted into a logical grid structure, i.e., decomposing the cross-row and cross-column cells into multiple logical grid cells. This effectively unifies the interpolation style of each cell and improves the regression accuracy of the interpolation map. Instead of using cells to generate interpolation maps before training, grids are used.
4 Experiments
4.1 Benchmark Datasets
We conducted experiments on three popular public benchmarks, including the ICDAR2013 Table Competition dataset, the Wired Table in the Wild (WTW) dataset, and the TableGraph-24k dataset to validate the effectiveness of TableCenterNet.
ICDAR-2013 [47] consists of 156 tables in PDF format from EU/US government websites, which contain cross-cells and various other styles. Because this paper focuses on reconstructing table from cells, an modified version of ICDAR-2013, the ICDAR-2013.c [48], is used, which provides detailed cell bounding boxes rather than word region boxes. It should be noted that the original dataset does not specify the training and test sets, we use 80% of the table images for training and others for test following the setting in [40, 43, 41, 5, 49].
WTW [7] consists of digital document images, scanned images, and images taken in the field, and contains a total of 10,970 training images and 3,611 test images collected from wild complex scenes. This dataset focuses on bordered tabular objects only and covers a wide range of challenging situations such as inclined tables, curved tables, occluded and blurred tables, extreme aspect ratio tables, overlaid tables, irregular tables, and multicolored tables. The labeling of the dataset includes both the physical and logical coordinates of the cells, with the physical coordinates represented by quadrilaterals.
TableGraph-24k [4] is a subset of TableGraph-350K, containing 20,000 training images, 2,000 validation images, and 2,000 test images. All images in this dataset are sourced from scholarly articles on the arXiv platform and include wired tables, wireless tables, and partial border tables.
4.2 Evaluation Metrics
Similar to previous studies [40, 4], we first evaluate the performance of our model in predicting the spatial locations of cells (physical coordinates) using precision, recall, and F1 scores under the IoU threshold of 0.5. For the logical locations of cells (row/column information), we employed the accuracy of logical location [50], the F1 score of adjacency relationships between cells [51, 47], and Tree-Edit-Distance-based Similarity (TEDS) [36] as evaluation metrics. The accuracy of logical location and TEDS directly reflect the correctness of the predicted structure, while the adjacency relationships only measure the quality of the intermediate results of the structure [5].
4.3 Implementation Details
All experiments were conducted in a PyTorch environment and executed on a workstation equipped with two GeForce RTX 3090 GPUs. In these experiments, we respectively adopted DLA-34 [52] and StartNet-s3 [53] as the backbone networks, and their weights were all initialized based on pre-trained models for the ImageNet [54] classification task. Within the CenterNet [10] framework’s regression network, we configured the number of hidden channels for each regression head to . Model optimization was performed using the Adam algorithm with a batch size of ( for TableGraph-24k). For both backbone networks, the models were trained for epochs, with the initial learning rate set to . The learning rate was decayed to and in the 140th and 180th epochs, respectively. To ensure scale invariance, we uniformly resized the input table images to a fixed size of pixels, and pixels for the TableGraph-24k dataset, while maintaining the aspect ratio of the images. After resizing, we also applied data augmentation and normalization to the input images to accelerate the convergence of the network.
Due to the limited size of the ICDAR-2013 dataset, which contains only 258 table images after preprocessing, we opted to perform rounds of fine-tuning on the model trained on WTW using the segmented ICDAR-2013 training set. During the fine-tuning phase, considering that most tables in the WTW dataset have outer margins, we extend the original image by 100 black pixels on all sides before scaling the image. The initial learning rate decays to and in the 70th and 90th epochs, respectively, while ensuring that the other training configurations remain unchanged. In the testing phase, we similarly expand the perimeter of the test image by black pixels before performing inference.
4.4 Results
To evaluate TableCenterNet’s ability to recognize table structures in different scenarios, we tested it on three benchmark datasets with different scenarios. The qualitative results in Fig. 6 show that our method is robust to complex structures, diverse colors, geometric transformations, and distorted or deformed tables in both photographic and scanning scenes. Furthermore, our method also demonstrates generalization to digital native scenarios, effectively handling tables across cells, with bounded cells, and without bounded cells.
Results on ICDAR-2013. As shown in Table 1, our proposed TableCenterNet method outperforms most previous methods in recognizing both the physical and logical structures of tables. Compared to the robust baseline model LORE [5], TableCenterNet-D enhances the physical coordinate F1 scores and the accuracy of the logical location by 0.6% and 4.3%, respectively. Similarly, TableCenterNet-S improves these two metrics by 0.3% and 1.3%, respectively. For comparison with Cycle-CenterNet [7], we used a model trained on the WTW dataset to test the wired tables in ICDAR-2013 [48]. The results indicate that our method exceeds Cycle-CenterNet [7] and SCAN [8] in terms of precision, recall, and F1 score metrics for adjacency. This shows that our method achieves competitive results in both wired and wireless scenarios.
Model | Backbone | Training Datasets | Physical Coordinates | Adjacency Relation | Acc | ||||
P | R | F1 | P | R | F1 | ||||
ReS2TIM [50] | - | ICDAR2013 | - | - | - | 73.4 | 74.7 | 74.0 | 17.4 |
TGRNet [4] | ResNet-50 | ICDAR2013 | 68.2 | 65.2 | 66.7 | - | - | - | 27.5 |
Cycle-CenterNet [7] | DLA-34 | WTW+ICDAR2019 | - | - | - | 95.5 | 88.3 | 91.7 | - |
TabStrNet [40] | ResNet-101 | SciTSR | - | - | - | 93.0 | 90.8 | 91.9 | - |
LGPMA [25] | ResNet-50 | ICDAR2013 | - | - | - | 96.7 | 99.1 | 97.9 | - |
Cycle-CenterNet [7] | DLA-34 | WTW | - | - | - | 97.5 | 98.4 | 98.0 | - |
TOD [45] | - | FinTabNet | - | - | - | 98.0 | 97.0 | 98.0 | - |
SCAN [8] | ResNet-50 | WTW | - | - | - | 98.2 | 98.1 | 98.2 | - |
FLAGNet [41] | ResNet-50 | ICDAR2013 | - | - | - | 97.9 | 99.3 | 98.6 | - |
NCGM [43] | ResNet-18 | ICDAR2013 | - | - | - | 98.4 | 99.3 | 98.8 | - |
LORE [5] | DLA-34 | ICDAR2013 | - | - | 97.2 | 99.2 | 98.6 | 98.9 | 86.8 |
TableCenterNet-D | DLA-34 | ICDAR2013 | 98.7 | 96.9 | 97.8 | 96.1 | 99.1 | 97.6 | 91.1 |
TableCenterNet-D | DLA-34 | WTW | 97.1 | 97.9 | 97.5 | 98.4 | 98.6 | 98.5 | 94.6 |
TableCenterNet-S | StartNet-s3 | ICDAR2013 | 98.2 | 96.8 | 97.5 | 93.7 | 97.5 | 95.6 | 88.1 |
TableCenterNet-S | StartNet-s3 | WTW | 97.1 | 98.2 | 97.6 | 98.4 | 98.6 | 98.5 | 94.1 |
Results on WTW. Considering that the previous methods used mainly two IoU thresholds, 0.5 and 0.9, to evaluate the prediction performance of physical coordinates, we conducted independent evaluations under these two distinct thresholds. The experimental results are presented in Table 2. The physical coordinate F1 scores of TableCenterNet-D and TableCenterNet-S are both improved by 0.9% compared to the optimal LORE [5] at an IoU threshold of 0.5, and the physical coordinate F1 scores at an IoU threshold of 0.9 demonstrated improvements of 13% and 12.6% over Cycle-CenterNet [7], respectively. For the evaluation of cell logical coordinates, TableCenterNet-D outperforms LORE [5] in both adjacency and the accuracy of logical location. In addition, TableCenterNet-D and TableCenterNet-S improve by 1% and 0.4% in TEDS compared to the optimal method [8], indicating that our approach also offers parsing advantages for complex tables in the wild.
Model | Backbone | Physical Coordinates | Adjacency Relation | Acc | TEDS | ||||
P | R | F1 | P | R | F1 | ||||
CascadeTabNet [55] | CascadeNet | - | - | - | 16.4 | 3.6 | 5.9 | - | 11.4 |
Split+Heuristic [27] | - | 3.2 | 3.6 | 3.4 | 25.7 | 29.9 | 27.6 | - | 26.0 |
TGRNet [4] | ResNet-50 | - | - | 64.7 | - | - | - | 24.3 | - |
FLAGNet [41] | - | - | - | - | 91.6 | 89.5 | 90.5 | - | - |
Cycle-CenterNet [7] | DLA-34 | 78.0 | 78.5 | 78.3 | 93.3 | 91.5 | 92.4 | - | 83.3 |
TSRFormer [31] | ResNet-18 | - | - | - | 93.7 | 93.2 | 93.4 | - | - |
NCGM [43] | ResNet-18 | - | - | - | 93.7 | 94.6 | 94.1 | - | - |
SCAN [8] | ResNet-50 | - | - | - | - | - | - | - | 90.7 |
TRACE [9] | ResNet-50 | 63.8 | 65.7 | 64.8 | 93.5 | 95.5 | 94.5 | - | - |
LORE [5] | DLA-34 | - | - | 96.4 | 94.5 | 95.9 | 95.1 | 82.9 | - |
TableCenterNet-D | DLA-34 | 98.1 | 96.5 | 97.3 | 94.8 | 97.2 | 96.0 | 83.0 | 91.7 |
TableCenterNet-D | DLA-34 | 91.9 | 90.8 | 91.3 | - | - | - | - | - |
TableCenterNet-S | StartNet-s3 | 98.1 | 96.5 | 97.3 | 94.3 | 96.9 | 95.6 | 81.5 | 91.1 |
TableCenterNet-S | StartNet-s3 | 91.4 | 90.4 | 90.9 | - | - | - | - | - |
Results on TableGraph-24k. To our knowledge, only two methods have been trained and tested on this data set for the precision of cell physical and logical coordinates: TGRNet [4] and LORE [5], respectively. In Table 3, we observe that TableCenterNet significantly outperforms previous methods. When the F1 scores for the physical coordinates of the cells are quite similar compared to the optimal method [5], TableCenterNet-D and TableCenterNet-S have improved the accuracy of logical location by 7.2% and 6.4%, respectively, achieving state-of-the-art performance. This demonstrates the substantial advantage of our methods in the digital native table image recognition scenario for scientific articles.
Model | Backbone | Physical Coordinates | Logical Location | ||||||
P | R | F1 | AccrowSt | AccrowEd | AcccolSt | AcccolEd | Acc | ||
TGRNet [4] | ResNet-50 | 91.6 | 89.5 | 90.6 | 91.7 | 91.6 | 91.9 | 92.3 | 83.2 |
LORE [5] | DLA-34 | - | - | 96.1 | - | - | - | - | 87.9 |
TableCenterNet-D | DLA-34 | 95.4 | 96.6 | 96.0 | 97.9 | 97.9 | 97.3 | 97.1 | 95.1 |
TableCenterNet-S | StarNet-s3 | 95.7 | 96.6 | 96.2 | 97.8 | 97.7 | 96.7 | 96.4 | 94.3 |
4.5 Ablation Study
In this subsection, we isolate the contributions of the key components in TableCenterNet and conduct ablation experiments on the WTW dataset. In all the experiments presented in Table 4, the differences in spatial location accuracy are minimal, so we focus on the effect of key components on logical location accuracy.
Effectiveness of cell boundaries and spans supervisions. Based on the results of Exp.1a and Exp.1b, we observe that the cell-span regression yields a performance improvement of +4.4% Acc. Comparing Exp.1a and Exp.1c, the cell span supervised loss leads to an even more substantial enhancement, with a performance improvement of +6.5% Acc. This is due to the fact that provides anchors for interpolation plot regression and strengthens the constraints on the range of interpolation between cells. From Exp.1d and Exp.1f, the cell boundary supervised loss also improves by +0.6%Acc due to the fact that it reduces the interpolation interference. According to the properties of interpolation map generation, the logical indices of cell boundaries are closer to the true values, while the logical indices of in-cell interpolation are approximate and less reliable. In Fig. 7, we visualize the accuracy of each logical location by heat map. From the three heat maps, we can find that cell spanning regression improves the accuracy of small logical indices. When the two supervision mechanisms are combined, the accuracy of large logical indices is also significantly improved.
Impact of different logical location regression methods. In Exp.2a and Exp.2b, we replace interpolation map regression with logical indices regression of the upper-left corner point and the four corner points respectively. Both of them use the optimal loss computation scheme, and the network structure remains consistent with TableCenterNet. Compared with Exp.2f, the performance of these two logical location regression methods decreases by 3.2%Acc and 1.2%Acc, respectively, indicating that the interpolation map regression can better model the dependencies and constraints between logical locations of different cells. When there is an offset in the spatial locations regressed by the methods in Exp.2a and Exp.2b, the logical locations are prone to errors, whereas interpolation maps can ensure the accuracy of logical indices as much as possible by rounding the interpolation results.
Exp. | Regression | Supervisions | Spatial Location | Logical Location | ||||||
Logical Indices | Cell Span | P | R | F1 | A-c | A-r | Acc | |||
1a | Interpolation map | ✗ | ✗ | ✗ | 98.1 | 96.4 | 97.3 | 86.4 | 84.8 | 75.1 |
1b | Interpolation map | ✓ | ✗ | ✗ | 98.1 | 96.4 | 97.2 | 89.0 | 87.4 | 79.5 |
1c | Interpolation map | ✗ | ✓ | ✗ | 98.2 | 96.5 | 97.4 | 90.7 | 88.3 | 81.6 |
1d | Interpolation map | ✓ | ✓ | ✗ | 98.1 | 96.5 | 97.3 | 91.0 | 88.8 | 82.4 |
1f | Interpolation map | ✓ | ✓ | ✓ | 98.1 | 96.5 | 97.3 | 91.4 | 89.1 | 83.0 |
2a | Upper-left corner point | ✓ | - | - | 97.8 | 96.4 | 97.1 | 89.6 | 87.0 | 79.8 |
2b | Four corner points | ✓ | ✓ | - | 97.8 | 96.5 | 97.1 | 90.9 | 88.2 | 81.8 |
4.6 Computational Analysis
We compare three measures of the computational efficiency of the models in Table 5, provided that the comprehensive performance is comparable. The input image size is , and both the comprehensive performance and the average inference time are computed based on the WTW [7] test set, where the comprehensive performance is referenced from [4], based on a combination of the F1 of the physical coordinates and the accuracy of logical locations. With the same backbone, TableCenterNet-D reduces the number of parameters by 27.3% compared to LORE, but increases the inference speed by 15.7 times. TableCenterNet-S further improves the computational efficiency by using the StarNet-s3 [53] backbone. Compared with TableCenterNet-D, the number of model parameters and FLOPs of TableCenterNet-S are reduced by 62.7% and 62.2%, respectively, which makes it possible to deploy on edge devices.
Model | Backbone | Fβ=0.5 | Parameters | FLOPs | Average Inference Time |
LORE | DLA-34 | 0.934 | 27.86 | 142.3 | 3788.4 |
TableCenterNet-D | DLA-34 | 0.941 | 16.82 | 133.1 | 227.4 |
TableCenterNet-S | StarNet-s3 | 0.937 | 6.27 | 50.3 | 214.8 |
5 Conclusion
In this paper, we present a one-stage end-to-end table structure parsing method, TableCenterNet, which uses a single model to simultaneously regress both the spatial and logical structures of tables. Unlike previous methods [4, 5] that focus on logical location regression, TableCenterNet directly regresses the distribution of the logical structure of each table at various locations within the input image only through convolution along with supervision of cell boundaries and spans, simplifying the model training process and accelerating model inference. Experimental results demonstrate that our approach achieves competitive outcomes across diverse scenarios, including the challenging table scenario, and achieves state-of-the-art performance on the TableGraph-24k dataset. In future work, we plan to collect challenging wireless tabular data from diverse scenarios to further train our models and provide a reference for improvement.
Acknowledgments
This work is supported by the Natural Science Foundation of China (No. 62366034).